# LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Automation Task Evaluation

### v1

@@ 3. LlamaTouch Design
Compared to previous evaluation approaches, LlamaTouch exhibits the following benefits.
$\bullet$ Practical on-device task execution ($\S$§3.1). Most of the previous evaluation methods are simulated on mobile UI interaction datasets. LlamaTouch, on the other side, enables mobile agent evaluation on real mobile devices, revealing their real capabilities in the wild.
$\bullet$ Fine-grained essential application state annotation ($\S$§3.2). By observing UI automation task execution only transfers essential UI elements within an application, LlamaTouch enables annotators to explicitly annotate essential application states that should be detected and matched during task execution.
$\bullet$ Faithful and scalable task evaluation ($\S$§3.3). LlamaTouch evaluates mobile agents by comparing task execution traces with annotated application states. By combining exact matching with fuzzy matching algorithms on different application states, LlamaTouch achieves faithful and scalable task evaluation.
3.1. On-device Task Execution
Evaluating mobile agents on pre-defined traces, as discussed in Section 2.2, depicts significant inaccuracy. LlamaTouch empowers on-device UI automation task execution to reveal the real capabilities of mobile agents. To simplify this process, LlamaTouch operates as a bridge between existing mobile agents and real mobile devices (e.g., smartphones, Android emulators, and cloud device farms). As shown in Figure 2, mobile agents execute tasks on LlamaTouch by taking the following processes. 1 Mobile agents request one task instruction from LlamaTouch. The instruction comes from one supported dataset by LlamaTouch, e.g., AITW. 2 With this task instruction, mobile agents will then request UI representation (e.g., screenshot, VH) from LlamaTouch, which will be 3 forwarded to mobile devices. 4 Taking the task instruction and UI representation as inputs, mobile agents will predict an action to be performed on the current UI, and dispatch predicted actions to LlamaTouch. 5 LlamaTouch forwards and executes the agent-predicted action to real mobile devices. The processes from 2 to 5 are repeatedly executed until mobile agents treat the task as completed. During task execution, all UI representations and corresponding actions are captured for further evaluation in $\S$§$3.$ These UI representations (1) pixel-level screenshots; (2) textual screen VHs; (3) activity names of the application at the foreground of each screen; and (4) actions performed on every screen. Essential system states such as the list of installed applications will also be dumped for faithful mobile agent evaluation, as we will illustrate in $\S$§3.2. Appendix A.1 details APIs provided by LlamaTouch for mobile agent integration.
Table 2. Primitives of essential states in LlamaTouch that are used in the annotation and evaluation processes.
@@ 3.1. On-device Task Execution
Fuzzy match UI state Screen info fuzzy<-1> Check if the contents on two screens are approximately identical.
@@
Check if the content of the target textbox is semantically similar
@@
A coarse-grained approach to determining if two UIs represent
@@
Check if the UI component is exactly identical to the UI component<n>,
or does not occur, in the ground-truth UI.
@@
Check if the target application named "app" has been successfully
installed/uninstalled.
@@
Check if two actions and their parameters are identical.
Figure 3. Annotated essential state for task “empty the shopping cart on bestbuy” in the last UI representation: two textboxes that should be exactly matched. The essential state represents the application state after task execution: “the shopping cart on bestbuy is empty”.
@@ 3.2. Essential Application State Annotation
Two major cases demonstrated in $\S$§2.2 make exact matching on pre-defined traces unfaithful. First, it demands that the agent-generated actions and their parameters match exactly with those in the dataset. Second, it takes a fixed UI interaction sequence provided by the dataset as the reference, therefore unable to evaluate alternative task execution paths.
@@
Insight: the task execution process transfers identifiable application states. The internal of the execution process of a UI automation task is changing application states that some of them can be explicitly represented by UI components. Even if the task execution paths differ, there are overlaps in the essential application states. We can utilize these overlapping application states to determine whether a task achieves some milestones or is completed. As the example shown in Figure 1(b), by identifying the intent of the task “open app Microsoft Excel (install if not already installed), go to login, …”, it contains two potential essential states for evaluation: (1) the application “Microsoft Excel” is opened; and (2) the application is located at the login page. Other actions, like detecting whether the application is installed, can be omitted during evaluation.
@@
To do so, essential states should be accurately identified and annotated. However, simply annotating application states at the whole UI representation level (e.g., a screenshot) and comparing screen-level similarity (Feiz et al., 2022), are far coarse-grained and may lead to inaccurate matching. For example, the screen contents of a web-shopping application could be subtly different due to slight movement or dynamically loaded contents across different executions. For accurate and efficient essential state annotation, LlamaTouch breaks the whole pixel-level UI representation (i.e., screenshot) to separate UI components. LlamaTouch simplifies the textual VH of each screen, which precisely expresses the attributes of every UI element, to extract important, visible UI components. These extracted UI components are combined with visually intuitive screenshots to provide precise overlayed bounding boxes and unique identifiers to annotators. Figure 3 shows an example of the VH-enhanced screenshot provided to annotators.
@@
Primitives. LlamaTouch incorporates a list of primitives for essential state annotation. These primitives are used to extract essential information on the screen. Annotators are instructed to clearly identify and use primitives to represent application states that are informative and deterministic to validate the task execution results. Table 2 comprehensively shows these primitives and their use cases. Overall, LlamaTouch incorporates six types of primitives and nine keywords for annotation. These primitives can be divided into three types according to the state type they represent.
@@
$\bullet$ UI state. The purpose of UI state annotation is to extract and compare whether two screens contain identical or similar information. As a screen may contain different types of components, LlamaTouch has four major primitives to cover all these UI components. (1) Screen info and (2) textbox are used to compare whether the whole screen or a dedicated textbox has similar contents. Screen info is primarily used for checking whether two UI representations are within the same screen (i.e., the same page of an application). Textbox primitive is used to annotate textboxes that may contain dynamic contents, such as a search box in a web-shopping application or the URL of a website. (3) The functionality of activity is like screen info: it can be used to approximately detect whether two UI representations are on the same screen of an application. (4) UI component is used to deterministically annotate the state of a UI component, including textbox, button, image, etc. For example, if the content of a textbox should be identical, it should be annotated with the exact keyword. States that require certain UI components should not occur on the screen should be annotated with the exclude keyword.
@@
$\bullet$ System state. LlamaTouch involves system state annotation mainly inspired by the tasks in a previously built dataset: AITW (Rawles et al., 2023). For example, a task “install app YouTube Kids” can be detected directly using shell commands (e.g., pm list packages), without involving redundant UI operations. Currently, LlamaTouch supports programmatically checking application installation. The keywords can be easily extended when other system states are required to be detected.
@@
$\bullet$ Action. Although the principle of annotation is to detect application state transfer during task execution. Sometimes, action on a specific UI representation is necessary to validate agent behaviour, especially when there is not enough state identifier shown in UI. LlamaTouch provides click and type keywords that covers most common actions.
@@
Case study. Essential state annotation can be divided into the following processes. First, for each UI representation, LlamaTouch overlayed all functional UI components with numeric markers to the screenshot as shown in Figure 3. This is done by extracting the precise metadata of each UI component from textual VH. Then, annotators will identify what are essential states that should be checked during evaluation to ensure the task is completed. LlamaTouch simplify this process by only requiring annotators to explicitly identify what UI element should be matched. For example, after emptying the shopping cart on bestbuy, the screen will display a textbox with the content “Your cart is empty”. This textbox should be treated as an essential state, which is highlighted in a bounding box with numeric id 13 in the last screen of Figure 3, as it represents the state after task completion. As we anticipate the content of the textbox should be exactly matched, the annotated keyword is exact<13>. The keyword exact<27> is used to validate whether two screens are both in the shopping cart of bestbuy. All annotations along with the UI representations and task descriptions construct the essential state-powered dataset. These essential states are used by LlamaTouch for faithful task evaluation.
$3.$ Faithful and Scalable Task Evaluation
@@
Annotated essential states ($\S$§3.2) and task execution traces captured on real devices ($\S$§3.1) are used for evaluation on LlamaTouch. LlamaTouch iterates essential states to determine whether a task execution trace goes through all annotated states sequentially. If so, this task is deemed completed. To achieve faithful evaluation, the most significant challenge lies in how to ensure the task execution trace matches the essential states. Task execution traces captured from real-world environments may contain dynamic screen content; it is vital to adapt the annotated essential states on a static dataset to dynamic, real-world contents, while achieving precise match on only critical information. To address this problem, LlamaTouch employs a multi-level state matching algorithm, which combines fuzzy match and exact match on both the entire screen and separated UI components, to ensure faithful evaluation. It first approximately matches two screens according to their UI representation and activity. Then, within two matched screens, it compares the annotated primitives that should be fuzzily or exactly matched.
@@
Approximate screen match is used to ensure two screens are on the same functionality page of an application, even under high screen content dynamics. This is the prior process for in-screen exact/fuzzy content match. To achieve this, LlamaTouch utilizes two annotation primitives proposed in $\S$§3.2: activity and screen info.
@@
1. Application activity match. Activity represents an entry point for users to interact with the application (and, 2024). This is an obvious identifier to indicate whether two screens are in the same application. Screens with distinct functionalities typically have unique activity names, even if they are within the same application. For example, the activity of the main setting page is com.android.settings.Settings, while the value is com.wifiadmin.settings.WifiSettingsActivity in the Wi-Fi setting page. Typically, exact application activity match acts as a foundational filtering process to identify whether two screens are on the same page. However, some specific application design philosophies may make different functional pages of an application contain the same activity name (sin, 2018). Under this circumstance, LlamaTouch involve screen info for accurate screen match.
@@
2. Fuzzy screen info match is utilized when the activity name cannot reveal different functional pages of an application. To better compare two screens, the primary thing is to only extract critical information from the screen. In summary, LlamaTouch simplifies the textual screen VH to a simple HTML representation as in prior work (Wen et al., 2023; Wang et al., 2023), while containing the types of every single UI component. LlamaTouch compares two simplified HTML representations of two screens with a cosine similarity of their sentence embeddings. If the result falls within a threshold, LlamaTouch treats two screens are matched. Two screens are similar when they are above a predefined threshold, e.g., 0.85 in our experiments. The fuzzy screen info match design helps LlamaTouch maintain faithfulness during evaluation when dealing with dynamic screen contents.
@@
1. Exact UI component match requires the annotated UI component should be identical in two screens to be compared, including all their attributes such as class, text, and selected in VH. This is especially useful for evaluating the content of a textbox, the status of a button (e.g., checked or not), or a selectable icon on the screen. Given a target UI component to be exactly matched, LlamaTouch will iterate nodes in VH of the matched screen, until one matched UI component is found. Exact UI component match fails if there is no matched node on available screens. In our dataset, exact UI component match contributes to 50.5% (688 out of 1,362) of annotated essential states.
@@
2. Fuzzy textbox match is crucial for comparing content inside a textbox on the screen, especially when the content may be slightly different. Semantically similar searching keywords with the same intent that complies with a specific task description should be matched, as they will lead to the same results. For example, Figure 1(b) shows searching “Microsoft Excel” and “Excel” in the Google Play Store both display the target application. LlamaTouch extracts the content of the annotated textbox, and then compares the text with nodes in the matched screen using the same approach as in fuzzy screen info match.
@@
3. Action match. Although the initiative of LlamaTouch is to detect state transfer during task execution. There are still cases for evaluating concrete actions performed on the screen, such as clicking one specific UI component, or typing the correct captcha. LlamaTouch directly compares actions and their parameters of the annotated ones with their counterparts on the matched screens.
@@
4. System state match is usually more efficient and accurate than merely comparing UI states for specific tasks that involve deterministic system states, e.g., installed applications. LlamaTouch currently supports checking whether an application is installed or not. Such system state is recorded during task execution in real-world mobile devices ($\S$§3.1). LlamaTouch will check whether the annotated system state is identical to that of the last screen in the task execution trace.
@@
Through the multi-level state matching algorithm, LlamaTouch achieves high evaluation accuracy on real-world task execution traces, while preserving the scalability of evaluating on static datasets. Experiments in $\S$§4 show the faithful evaluation of LlamaTouch.

### v2

@@ 3. LlamaTouch Design
Figure 2 shows the workflow of using LlamaTouch for on-device task execution and trace evaluation, using the well-constructed LlamaTouch dataset. Compared to previous evaluation approaches, LlamaTouch exhibits the following benefits.
$\bullet$ Practical on-device mobile UI task execution (§$\S$3.1). Most previous evaluation methods are simulated on mobile UI interaction datasets. LlamaTouch, on the other hand, enables mobile agents to operate in realistic mobile environments for UI task automation, revealing their true capabilities in real-world scenarios.
$\bullet$ Fine-grained essential application state annotation (§$\S$3.2). By observing UI automation task execution transfers states of essential UI elements within an application, LlamaTouch enables annotators to explicitly annotate essential application states that should be detected and matched for task completion. This approach avoids the previous static evaluation methods’ focus on the determinism of traces, reducing the probability of false negatives in evaluation.
$\bullet$ Faithful and scalable task evaluation (§$\S$3.3). LlamaTouch evaluates the performance of mobile agents by comparing their task execution traces captured in real-world mobile environments with annotated essential states. By combining exact and fuzzy matching algorithms on different application states, LlamaTouch achieves faithful task evaluation without losing scalability.
3.1. On-device Mobile UI Task Execution
Table 2. A set of primitives used in essential state annotation and trace evaluation.
@@ 3.1. On-device Mobile UI Task Execution
Fuzzy match UI state Screen info fuzzy<-1> Verify if the contents on two screens are approximately identical.
@@
Verify if the content of the target textbox is semantically similar
@@
A coarse-grained approach to determine whether two UIs represent
@@
Verify if the UI component is exactly identical to the UI component<n>,
or if it does not occur, in the ground-truth UI.
@@
Verify if the target application named "app" has been successfully
installed or uninstalled.
@@
Verify if two actions and their parameters are identical.
Evaluating mobile agents on predefined, deterministic traces, as discussed in Section 2.2, depicts significant inaccuracy. LlamaTouch empowers on-device mobile UI task execution to reveal the real capabilities of mobile agents. To simplify this process, we propose AgentEnv, a bridge between existing mobile agents and realistic mobile environments (e.g., smartphones, Android emulators, and cloud device farms). With AgentEnv, mobile agents execute tasks on real-world mobile environments by following the processes shown in Figure 3. 1 Mobile agents request a task instruction from AgentEnv. The instruction comes from the LlamaTouch dataset. 2 With the task instruction, mobile agents request UI representations (e.g., screenshot, VH) from AgentEnv, which are then 3 forwarded to mobile devices. 4 Taking the task instruction and UI representations as inputs, mobile agents predict an action to be performed on the current UI and dispatch the predicted action to AgentEnv. 5 AgentEnv forwards and executes the agent-predicted action to mobile environments. The processes from 2 to 5 are repeated until mobile agents consider the task completed. During task execution, all UI representations and corresponding actions are captured as task execution traces for further evaluation in §$\S$3.3. These UI representations include (1) pixel-level screenshots; (2) textual screen VHs; (3) activity names of the application in the foreground of each screen; and (4) actions performed on each screen. Essential system states, such as the list of installed applications, will also be recorded for faithful mobile agent evaluation. The left-hand side of Figure 2 demonstrates the code implementation of the interaction between mobile agents and AgentEnv. Appendix A.1 presents details of the APIs provided by LlamaTouch for mobile agent integration.
Figure 3. Interaction between mobile agents, AgentEnv, and real-world mobile environment in the on-device mobile UI task execution process.
@@ 3.2. Essential Application State Annotation
Two major cases demonstrated in §$\S$2.2 highlight why exact action match on predefined action sequences is unfaithful. First, it demands that the agent-generated actions and their parameters match exactly with those in the dataset. Second, it takes a fixed UI interaction sequence provided by the dataset as the reference, making it unable to evaluate alternative task execution paths.
@@
Insight: The task execution process transfers identifiable application states. During the execution of a UI automation task, the application states change, and some of these states can be explicitly represented by UI components. Even if the task execution paths differ, there are overlaps in the essential application states. We can utilize these overlapping application states to determine whether a task achieves some milestones or is completed. As the example shown in Figure 1, by identifying the intent of the task “open app Microsoft Excel (install if not already installed), go to login, …”, it contains two potential essential states for evaluation: (1) the application “Microsoft Excel” is opened; and (2) the application is located at the login page. Other actions, like detecting whether the application is installed, can be omitted during evaluation.
@@
To achieve this, essential states should be accurately identified and annotated. However, simply annotating application states at the whole UI representation level (e.g., an entire screenshot) and comparing screen-level similarity (Feiz et al., 2022) is too coarse-grained and may lead to inaccurate matching. For example, the screen contents of a web-shopping application could be subtly different due to nondeterministic swiping gestures or dynamically loaded contents across different executions. For accurate and efficient essential state annotation, LlamaTouch breaks the whole pixel-level UI representation into separate UI components. This is achieved by simplifying the textual VH of each screen, which precisely expresses the attributes of every UI element, to extract important, visible UI components. These extracted UI components are combined with visually intuitive screenshots to provide precise overlayed bounding boxes and unique identifiers to annotators. Figure 4 shows an example of VH-enhanced screenshots in a task provided to annotators.
@@
Annotation primitives. LlamaTouch incorporates a list of primitives for essential state annotation. These primitives represent essential information on the screen and indicate how this information should be matched. Annotators are responsible for clearly identifying and using primitives to represent application states that are informative and deterministic for validating task execution results. Table 2 comprehensively shows these primitives and their use cases. Currently, LlamaTouch incorporates six types of primitives and nine keywords for annotation. These primitives can be divided into three types according to the application state type they represent.
Figure 4. Annotated essential states for the task “empty the shopping cart on bestbuy” in the last UI representation: two textboxes with the exact keyword. The essential states represent the application state after task execution: “the shopping cart on bestbuy is empty”.
@@
$\bullet$ UI state. The purpose of UI state annotation is to extract and compare whether two screens contain identical or similar information. As a screen may contain different types of components, LlamaTouch uses four major primitives to cover the comparison logic of all these UI components. (1) Screen info and (2) textbox are used to compare whether the whole screen or a dedicated textbox has similar contents. Screen info is primarily used for checking whether two UI representations are within the same screen (i.e., the same page of an application). Textbox primitive is used to annotate textboxes that may contain dynamic contents, such as a search box in a web-shopping application or the URL of a website. (3) The functionality of activity is like screen info: it can be used to approximately detect whether two UI representations are on the same screen of an application. (4) UI component is used to deterministically annotate the state of a UI component, including textbox, button, image, etc. For example, if the content of a textbox should be matched, this textbox should be annotated with the exact keyword. States that require certain UI components not to occur on the screen can be annotated with the exclude keyword.
@@
$\bullet$ System state. System state annotation in LlamaTouch is primarily inspired by tasks from the previously built dataset: AITW (Rawles et al., 2023). For example, a task like “install app YouTube Kids” can be detected directly using shell commands to access system states (e.g., pm list packages), without involving complex UI states. Currently, LlamaTouch supports programmatically checking application installation status. The keywords can be easily extended and customized to detect other system or application states.
@@
$\bullet$ Action. Although the principle of annotation is to detect application state transfer during task execution. Sometimes, action on a specific UI representation is necessary to validate agent behavior, especially when there are not enough state identifiers shown in the UI. LlamaTouch provides click and type keywords that cover the most common actions.
@@
Case study. Essential state annotation can be divided into the following processes. First, for each UI representation, LlamaTouch overlays all functional UI components with numeric markers on the screenshot as shown in Figure 4. This is done by extracting the precise metadata of each UI component from textual VH. Annotators will then identify the essential states that should be checked during evaluation to ensure the task is completed. LlamaTouch simplifies this process by only requiring annotators to explicitly identify what UI element should be matched and what annotation primitive should be used. For example, after emptying the shopping cart on bestbuy, the screen will display a textbox with the content “Your cart is empty”. This textbox should be treated as an essential state, which is highlighted in a bounding box with numeric ID 13 in the last screen of Figure 4, as it represents the state after task completion. As we anticipate the content of the textbox should be exactly matched, the annotated keyword is exact<13>. The keyword exact<27> is used to validate whether two screens are both in the shopping cart of bestbuy. All annotations, along with the UI representations and task descriptions, construct the essential state-powered dataset. These essential states are used by LlamaTouch for faithful task evaluation.
3.3. Faithful and Scalable Task Evaluation
@@
Annotated essential states (§$\S$3.2) and captured task execution traces (§$\S$3.1) are jointly used for evaluation in LlamaTouch. LlamaTouch iterates through essential states to determine whether a task execution trace sequentially matches all annotated states. If so, the task is deemed completed. To achieve faithful evaluation, the most significant challenge lies in how to ensure the task execution trace matches the essential states. Task execution traces captured from real-world mobile environments may contain dynamic screen content; it is vital to adapt the annotated essential states from a static dataset to varying screens captured from the real world, while achieving precise matches on only critical information. To address this problem, LlamaTouch employs a multi-level state matching algorithm, which combines fuzzy match and exact match on both the entire screen and separated UI components to ensure faithful evaluation. The algorithm first approximately matches two screens according to their UI representations and activities. Then, within two matched screens, it compares each annotated essential state with iterated UI components in the target screen based on their annotated primitives. A task is considered completed only when all annotated essential states have matching counterparts. The matching philosophy of all annotation primitives is as follows.
@@
Approximate screen match is used to ensure two screens are on the same functionality page of an application, even under high screen content dynamics. This is the preliminary process for in-screen exact/fuzzy content match. To achieve this, LlamaTouch utilizes two annotation primitives proposed in §$\S$3.2: activity and screen info.
@@
1. Application activity match. Activity represents an entry point for users to interact with the application (and, 2024). This is a clear identifier to indicate whether two screens are in the same application. Screens with distinct functionalities typically have unique activity names, even if they are within the same application. For example, the activity of the main settings page is “com.android.settings.Settings”, while the Wi-Fi settings page has the activity “com.wifiadmin.settings. WifiSettingsActivity”. Typically, exact application activity match acts as a foundational filtering process to identify whether two screens are on the same page. However, some specific application design philosophies may cause different functional pages of an application to contain the same activity name (sin, 2018). Under these circumstances, LlamaTouch involves screen info for accurate screen match.
@@
2. Fuzzy screen info match is utilized when the activity name cannot differentiate between functional pages of an application. To better compare two screens, it is crucial to extract only the critical information from the screen. In summary, LlamaTouch simplifies the textual screen VH to a simple HTML representation, as in prior work (Wen et al., 2024; Wang et al., 2023), while preserving the types of every UI component. LlamaTouch compares two simplified HTML representations of the screens using the cosine similarity of their sentence embeddings. Two screens are deemed similar when their cosine similarity exceeds a predefined threshold, e.g., 0.85 in our experiments. The fuzzy screen info match design helps LlamaTouch maintain faithfulness during evaluation when dealing with dynamic screen contents.
Figure 5. Workflow for constructing the LlamaTouch dataset. Generally, the workflow has three independent processes. (1) Generate task descriptions by sampling from previous datasets or constructing new ones through humans or LLMs. (2) Record UI interaction traces according to task descriptions. (3) Annotate essential states atop UI interaction traces.
@@
1. Exact UI component match requires that an annotated UI component be identical in two screens being compared, including all their attributes such as class, text, and selected in VH. This is especially useful for evaluating the content of a textbox, the status of a button (e.g., checked or not), or a selectable icon on the screen. Given a target UI component to be exactly matched, LlamaTouch will iterate through nodes in the VH of the matched screen until a matching UI component is found. Exact UI component match fails if no matching node is found on available screens. In our dataset, exact UI component match accounts for 51% (698 out of 1,379) of annotated essential states.
@@
2. Fuzzy textbox match is crucial for comparing the content inside a textbox on the screen, especially when the content may be slightly different. Semantically similar search keywords with the same intent that comply with a specific task description should be matched, as they will lead to the same results. For example, Figure 1 shows searching for “Microsoft Excel” and “Excel” in the Google Play Store both display the target application. LlamaTouch extracts the content of the annotated textbox and then compares the text with nodes in the matched screen using the same approach as in fuzzy screen info match.
@@
3. Action match. Although the initiative of LlamaTouch is to detect state transfer during task execution, there are still cases for evaluating concrete actions performed on the screen, such as clicking a specific UI component or typing the correct captcha. LlamaTouch directly compares the actions and their parameters of the annotated actions with their counterparts on the matched screens. Different from prior studies (Rawles et al., 2023), LlamaTouch uses XPath (xpa, 2017) of target UI components extracted from screen VHs as parameters for click actions. This approach provides more tolerance in the action space compared to using precise coordinates.
@@
4. System state match is usually more efficient and accurate than merely comparing UI states for specific tasks that involve deterministic system states, e.g., installed applications. LlamaTouch currently supports checking whether an application is installed or not. Such system states are recorded during task execution on real-world mobile devices (§$\S$3.1). LlamaTouch will check whether the annotated system state is identical to that of the last screen in the task execution trace, which is also recorded by AgentEnv.
@@
Through the multi-level state matching algorithm, LlamaTouch achieves high evaluation accuracy on real-world task execution traces, while preserving the scalability of evaluating on static datasets. The above evaluation logic is well encapsulated into the LlamaTouch evaluator. To use it for evaluation, an agent only needs to define how to load the task execution traces for each task, as shown in the right side of Figure 2, The evaluator will automatically conduct the evaluation and report metrics such as task completion rate. Experiments in §$\S$5 show the faithful evaluation of LlamaTouch.
4. LlamaTouch Dataset
4.1. Dataset Construction
The LlamaTouch dataset consists of a combination of tasks from the previous AITW dataset (Rawles et al., 2023) and self-generated tasks that involve diverse categories and popular applications. The inclusion of new tasks in currently popular applications, which have not been covered in previous literature, aims to properly assess the generality and real capabilities of mobile agents. Figure 5 demonstrates the workflow for constructing the dataset, where each data sample undergoes the following processes.
Generate task descriptions. Task descriptions are generated by both humans and LLMs using app metadata (e.g., app names, categories, descriptions) from popular apps in the Google Play Store. For tasks in the AITW dataset, we sample parts of them after deduplicating similar task descriptions. The sampled task descriptions are then validated by humans to avoid duplication, infeasibility, and high complexities that may far exceed the capabilities of human and mobile agents. After this process, we sampled 102 tasks from the AITW dataset among 26 unique apps and 394 newly generated tasks among 46 unique apps. The generated descriptions cover a variety of application categories such as utilities (e.g., Zoom, Expedia), social media (e.g., Discord, Instagram), and web shopping (e.g., Walmart, Amazon). The left-hand side of Figure 5 shows the proportion of these categories.
We then employ six human annotators, all of whom are authors of this study and experts in smartphone usage, to generate data samples in LlamaTouch through the following two independent annotation stages. The image on the lower side of Figure 5 illustrates the outputs of the human annotation process.
Record UI interaction traces. The validated task descriptions are used as guidance for recording UI interaction traces. Tasks sampled from the AITW dataset are also required to go through this process as they lack view hierarchies. Given a task description, human annotators interact with mobile apps through our developed UI Interaction Recorder. This tool is built on top of mobile emulators or realistic smartphones and displays the graphical user interface to users, allowing them to operate it like normal smartphones. Specifically, human annotators are asked to complete a task in the simplest manner and to avoid redundant operations, as in (Rawles et al., 2023). As shown in Figure 5, the task “Empty the shopping cart on bestbuy” requires five continuous click actions to complete. The recorder captures actions, VHs, activities, and screenshots, which collectively form the UI trajectory.
Annotate essential states. The recorded UI interaction traces, along with task descriptions, are then processed by human annotators to identify essential states. To help annotators better understand the application state transformation and simplify the annotation process, we developed an essential state annotation system. This system displays the entire UI interaction trace, with potentially significant UI components shown with numeric indices in each screenshot. Annotators are asked to identify the most significant, identifiable states that represent key milestones during task completion. They annotate these states (their numeric indices) with the proper essential state primitives we proposed in §$\S$3.2. These essential states, together with recorded UI interaction traces, form the final LlamaTouch dataset.
During the annotation process, to ensure data reliability, anything that a single human annotator cannot decide on is annotated based on a consensus reached by three or more annotators. Overall, this dataset includes 496 tasks, covering 57 unique Android applications with diverse task complexities.
4.2. Dataset Statistics
In this section, we present statistics of the LlamaTouch dataset. Table 3 quantifies task complexities by showing the average steps (actions) required to complete a task. Tasks from AITW (Rawles et al., 2023) are slightly more complex than those generated by LlamaTouch, with average steps of 7.35 versus 5.67, respectively. Overall, the average number of steps to complete a task is nearly 7, with task complexities ranging from 2 to 42 steps.
Table 3. LlamaTouch dataset statistics and task complexities measured by the average steps (actions) to complete a task.
Category # Task # Apps Avg. Steps
AITW (Rawles et al., 2023) 102 26 7.35 (2-19)
Generated 394 46 5.67 (3-42)
Total 496 57 7.01 (2-42)
We further analyze action types in the ground-truth dataset. Table 4 shows the statistics of actions contributing to the dataset. In summary, all 496 tasks involve click actions. Out of these, 292 tasks involve swipe actions, and 147 tasks involve input actions. A few tasks also use press home and back for navigation. By dividing the total number of actions by the number of tasks involving these actions, we observe an average of more than 4 click actions per data sample. The average number of all other actions is less than 2.
Table 4. Action statistics in LlamaTouch dataset.
Action # Action # Tasks W/ Action Mean/Stddev
Click 2,192 496 4.42/2.51
Swipe 376 292 1.29/0.9
Input Text 173 147 1.18/0.55
Press Home 44 43 1.02/0.15
Press Back 6 5 1.2/0.45
Table 5. Statistics of annotated essential states.
Type Essential State (ES) # ES # Tasks W/ ES Mean/Stddev
Exact Match UI Component 698 418 1.67/0.92
Activity 490 442 1.11/0.32
Action: Click 98 93 1.05/0.23
Action: Type 1 1 1/0
System: Install 7 7 1/0
System: Uninstall 3 3 1/0
Fuzzy Match UI Component 51 44 1.16/0.43
Screen Info 31 31 1/0
Statistics in Table 5 detail the annotated essential states. We find that the essential states of 441 tasks can be presented on only one UI representation, indicating that most tasks check only the final states after task completion. There are 53 tasks and 2 tasks with essential states presented on two and three UI representations, respectively. Among the annotated essential states, exact UI component match accounts for 51% of all essential states, followed by exact activity match at 36%. More than 84% of tasks (418 out of 496) have at least one of these two types of essential states. Fuzzy match is also important for the evaluation process: 44 tasks have 51 fuzzy UI component match in total; fuzzy screen info match occur in 31 tasks. In summary, all types of essential states construct the dataset, which significantly contributes to the faithful evaluation of mobile UI task automation, as we will show in the following section.